OpenStudy (anonymous):

simplify the expression b + 2/9 = 1 3/5 will give medal to best answer.

OpenStudy (anonymous):

\[b + \frac{ 2 }{ 9 } = 1 \frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

turn 1 3/5 into an improper fraction 1 3/5 = (1 x 5 + 3)/5 = 8/5 now we have : b + 2/9 = 8/5 (subtract 2/9 from both sides) b = 8/5 - 2/9 (common denominator is 45) b = 72/45 - 10/45 b = 62/45 or 1 17/45

OpenStudy (anonymous):

ok, thank yo!

OpenStudy (anonymous):

you* lol

OpenStudy (anonymous):

or b + 2/9 = 8/5 (multiply by the LCD, which is 45) 45b + 10 = 72 45b = 72 - 10 45b = 62 b = 62/45

OpenStudy (anonymous):

glad to help :)